5.7 An investment will pay $50 at the end of each of the next 3 years,...

5.7

An investment will pay $50 at the end of each of the next 3 years, $250 at the end of Year 4, $400 at the end of Year 5, and $600 at the end of Year 6. If other investments of equal risk earn 7% annually, what is its present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present value: $  

Future value: $  
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=50/1.07+50/1.07^2+50/1.07^3+250/1.07^4+400/1.07^5+600/1.07^6

=$1006.94(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value=1006.94*(1.07)^6

=$1511.14(Approx).